OpenStudy (anonymous):

A paint can can hold 2.5 litres, work out the radius of the can if the length is 17cm.

OpenStudy (compassionate):

Do you know the formula for radius?

OpenStudy (cgreenwade2000):

Do you know the formula for volume of a cylinder?*

OpenStudy (anonymous):

Volume of a cylinder = pi * r ^2 * h 2.5 = 3.14 * r^2 * 17 2.5 = r^2 * 53.38 r^2 = 53.38/2.5 = 21.352 r = sqrt(21.352) = 4.52 so, radius = 4.52.

OpenStudy (cgreenwade2000):

@kelly226 We're all really glad that you know how to do the math for this problem. But can you let the asker work it out themselves?

OpenStudy (anonymous):

sry! yes